E15-16 (Computation of Retained Earnings) The following information has been taken from the ledger accounts of Isaac Stern Corporation.

Total income since incorporation                              $317,000                    

Total cash dividends paid                                         60,000

Total value of stock dividends distributed                30,000

Gains on treasury stock transactions                       18,000

Unamortized discount on bonds payable                   32,000

Instructions

Determine the current balance of retained earnings.
Answer #1
Concepts and reason

The concept used to compute the retained earnings.

Financial Statement: The statements that are prepared for a specific period of time, comprising the financial information of the organization are known as financial statements. It includes statement of income which shows the profitability of the business, balance sheet which shows the monetary position with regards to assets, liabilities and capital, and cash flow statement which represents flow of cash for the accounting period.

Computation of Retained earnings: A computation statement that shows the changes in retained earnings during a financial year is the statement of retained earnings. It basically shows how the profits of a particular period are divided into distribution to shareholders and that retained to be used for the growth of the business.

Fundamentals

Retained Earnings: Retained earnings of a company is that portion of the company’s earnings that is kept by the company rather than distributing it as dividends.

Income: Income refers to the excess of revenue over the expenses in the financial year. It is the amount received from work or capital or land.

Dividend: Dividend is that portion of a company’s profits which are to be distributed to the shareholders. It is decided by the board of directors of the company.

Treasury Stock: Also known as reacquired shares, treasury stocks are those shares that were once outstanding shares of the company but are now repurchased by the company and further decommissioned.

Bond: An instrument used by a company to raise its capital by way of debt is called a bond. It is one of the company’s long-term liabilities. It is a type of loan which the company is obliged to repay to the holder of the bond after a definite time period.

Unamortized Discount: It refers to the discount amount on a debt has not written off yet in the income statement.

Computation of the total dividend distributed amount during the year:

Working Notes:

Computation of total dividend distributed:

TotalDividendDistributed=CashDividend+StockDividend=$60,000+$30,000=$90,000\begin{array}{c}\\{\rm{Total Dividend Distributed}} = \;{\rm{Cash Dividend}} + \;{\rm{Stock Dividend}}\\\\ = \;\$ 60,000 + \;\$ 30,000\\\\ = \;\$ 90,000\\\end{array}

Computation of Retained Earnings for I Corporation:

Working Note:

RetainedEarnings=(TotalIncomesinceIncorporationTotalCashDividendPaidValueofStockDividenddistributed)=($317,000$60,000$30,000)=$227,000\begin{array}{c}\\{\rm{Retained Earnings}} = \;\left( \begin{array}{l}\\{\rm{Total Income since Incorporation}} - \\\\Total\;Cash{\rm{Dividend}}\;{\rm{Paid}} - \\\\{\rm{Value of Stock Dividend distributed}}\\\end{array} \right)\\\\ = \left( {\$ 317,000 - \;\$ 60,000 - \;\$ 30,000} \right)\\\\ = \;\$ 227,000\\\end{array}

Ans:

Amount of current balance of retained earnings at the end of the year is $227,000.